Mason Company has two manufacturing departments—Machining and Assembly. The company considers all of its manufacturing overhead costs to be fixed costs. It provided the following estimates at the beginning of the year as well as the following information with respect to Jobs A and B:

Estimated Data Machining Assembly Total
Manufacturing overhead $ 1,352,000 $ 169,000 $ 1,521,000
Direct labor hours 13,000 104,000 117,000
Machine hours 104,000 8,000 112,000
Job A Machining Assembly Total
Direct labor hours 5 10 15
Machine hours 11 2 13
Job B Machining Assembly Total
Direct labor hours 4 5 9
Machine hours 12 3 15

Required:

1. If Mason Company uses a plantwide predetermined overhead rate with direct labor-hours as the allocation base, how much manufacturing overhead cost would be applied to Job A? Job B? (Round your answers to the nearest whole dollar amount.)

2. Assume that Mason Company uses departmental predetermined overhead rates. The Machining Department is allocated based on machine-hours and the Assembly Department is allocated based on direct labor-hours. How much manufacturing overhead cost would be applied to Job A? Job B? (Round your intermediate calculations and final answers to 2 decimal places.)
Answer #1

1) Overhead rate = 1521000/117000 = 13 per labor hour

Job A Job B
Manufacturing overhead cost applied 15*13 = 195 9*13 = 117

2) Machining overhead rate = 1352000/104000 = 13 per MH

Assembly overhead rate = 169000/104000 = 1.625 per DLH

Job A Job B
Machining overhead 11*13 = 143 12*13 = 156
Assembly overhead 10*1.625 = 16.25 8.125
Total overhead 159.25 164.13